# A Survey on Self-play Methods in Reinforcement Learning

### v3

@@ 3.5.5 Estimation Variants of CFR
Although these CFR variants advance the field, they can’t directly solve large imperfect-information extensive-form games due to their reliance on tabular representations. The typical approach involves abstracting the original game, applying CFR to the abstracted game, and translating strategies back to the original. This abstraction is game-specific and relies heavily on domain knowledge. Additionally, smaller abstractions often yield suboptimal results. Given these challenges, Waugh et al. (2015) introduces Regression CFR (RCFR), which employs a shared regressor $\varphi(I,a)$ to estimate counterfactual regrets. Nevertheless, using regression trees as the regressor limits RCFR’s applicability to miniature games, and the necessity for manually crafted features remains a drawback. After advantage-based regret minimization (ARM) (Jin et al., 2018) merges CFR with deep RL in single-agent scenarios, a growing body of research has focused on applying CFR in conjunction with neural networks to multi-agent scenarios. Double Neural CFR (Li et al., 2018) utilizes two neural networks: one for estimating counterfactual regrets and another for approximating the average strategy. In a similar vein, Deep CFR (Brown et al., 2019) leverages an advantage network $V(I,a|\theta_{p})$ to estimate counterfactual regrets with each player having a distinct hyperparameter $\theta_{p}$ and employs $\pi(I,a|\theta_{\pi})$ for strategy estimation after the training process of the advantage network. Since these two networks are trained in sequence rather than concurrently, the strategy for each intermediate iteration remains conditioned on the output of the advantage network: $h(i)=V(I,a|\theta_{p})$. Despite similarities, Deep CFR distinguishes itself from Double Neural CFR through its data collection and proven effectiveness in larger-scale poker games. Single Deep CFR (SD-CFR) (Steinberger, 2019) demonstrates training an average strategy network is unnecessary, with only an advantage network required. Building on the foundation of SD-CFR, DREAM (Steinberger et al., 2020) utilizes a learned baseline to maintain low variance in a model-free setting when only one action is sampled at each decision point. Moreover, advantage regret-matching actor-critic (ARMAC) (Gruslys et al., 2020) incorporates the thought of retrospective policy improvement.
@@
Table 1: Overview of representative self-play algorithms within our framework.
Algorithms MSS h(i) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{i}^{h}$ Suitable Game
Vanilla Self-play Equ. (5) $\emptyset$ Traditional Self-play $E=1$ Lazy General Stage Game
FSP Equ. (6)
@@
DO NE-based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Lazy General Stage Game
@@
FTW Equ. (10) $\emptyset$ Ongoing-training-based $E>1$ Immediate $\pi_{i}^{h}\leftarrow\pi_{i}(\cdot|h(i))$ Stage Game
NeuPL General $\sigma_{i}$
Simplex-NeuPL General $\sigma_{i}$
Vanilla CFR Equ. (5) $R^{i}(I,a)$ Regret-minimization-based $E=1$ Lazy $\pi_{i}^{h}\leftarrow\pi_{i-1}(\cdot|h(i-1))$ Repeated Game
CFR+ $R^{+,i}(I,a)$
CFR-D $R^{i}(I_{*},a)$
@@ 3.6 Reassessment of the Framework
After introducing these four categories of self-play algorithms, we will further compare them in this section, explain the rationale behind categorizing the algorithms, and summarize the representative algorithms in Table 1. Moreover, we will illustrate the differences between our framework and other frameworks to demonstrate why our proposed framework is more general.
3.6.1 Categorization Rationale
@@
Traditional self-play algorithms and the PSRO series share many similarities. Initially, they require only one randomly initiated policy, and the policy population expands as training progresses. Therefore, in our framework, we use lazy initialization to initialize the policy population and set $E=1$ for these two categories. The interaction matrix is typically a lower triangular matrix (Corollary 1 and Corollary 2). The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSes to handle tasks with intricate game-theoretical requirements. For example, $\alpha$-PSRO (Muller et al., 2020) specifically utilizes an $\alpha$-rank-based MSS to tackle multi-player general-sum games. In other tasks, traditional self-play algorithms are more commonly used to reduce the computational cost associated with complex MSSes in the PSRO series.
@@
Unlike the two previously mentioned categories, the ongoing-training-based series adopts a different paradigm. Instead of gradually expanding the policy population and relying on newer policies to be stronger, this approach strengthens all policies simultaneously at each epoch. This method alleviates issues such as early truncation and repeated skill relearning that occur in the above two categories. To integrate this category into our framework, immediate initialization is used for the policy population, and $\pi_{i}(\cdot|h(i))$ is utilized to initialize $\pi_{i}^{h}$ to ensure that policy updates are self-referential. Also, the interaction matrix is generally not a lower triangular matrix (Corollary 3).
Lastly, the regret-minimization-based series focuses on overall performance over time rather than on individual episodes, making it particularly suitable for repeated games. For example, Texas Hold’em is a classic repeated game where players adjust their strategies based on past interactions and use tactics involving deception and bluffing. The main goal of this training process is to update the regrets associated with different strategies, which in turn demands significant storage resources. Our framework uses $h(i)$ to store this information. Since the policies are determined by $h(i)$, only the most recent policy is relevant. Therefore, the interaction matrix is a unit lower shift matrix (Corollary 4). We also do not need actually to initialize the whole policy population and only need to use $\pi_{i-1}(\cdot|h(i-1))$ to initialize $\pi_{i}^{h}$ in the training process.
3.6.2 Comparing with Existing Frameworks
@@
Our framework is built upon PSRO (Lanctot et al., 2017) and NeuPL (Liu et al., 2022c). Here, we outline the differences between our framework and these existing frameworks. The primary distinction between our framework and PSRO is the use of an interaction matrix $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$ to represent the opponent sampling strategy, allowing for the integration of more complex competitive dynamics. Moreover, in our framework, $\sigma_{i}$ denotes the opponent sampling strategy, which specifies how to sample opponents’ policies against policy $i$ rather than being the meta-strategy of policy $i$. Additionally, our framework incorporates a policy condition function $h(i)$, making it more general than NeuPL, where policies are conditioned on $\sigma_{i}$. This enhancement gives our framework greater expressiveness. Furthermore, we describe how to compute the oracle (Line 5 in Alg. 1) in three different ways (Alg. 2, Alg. 3 and Alg. 4) to provide a clearer understanding. Also, to the best of our knowledge, our framework is the first self-play framework to integrate the regret-minimization-based series, which is a significant self-play paradigm.
4 Empirical Analysis
@@ 4.1 Board Games
Board games, the majority of which are perfect information games, were previously revolutionized by the introduction of two essential techniques: position evaluation and Monte Carlo tree search (MCTS) (Coulom, 2006; Kocsis & Szepesvári, 2006). These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess (Campbell et al., 2002), checkers (Schaeffer et al., 1992), othello (Buro, 1999), backgammon (Tesauro & Galperin, 1996), and Scrabble (Sheppard, 2002). In contrast, the application of these techniques to the game of Go with an estimated $2.1\times 10^{170}$ legal board configurations, only enabled performance at the amateur level (Bouzy & Helmstetter, 2004; Coulom, 2007; Baudiš & Gailly, 2011; Enzenberger et al., 2010; Gelly & Silver, 2007). In light of this, our discussion will specifically focus on the game of Go to illustrate the application of self-play. In addition to Go, we will explore Stratego, a board game with imperfect information, in contrast to most board games that only involve perfect information.
4.1.1 Go
@@
Go is an ancient board game, which is played on a grid of 19x19 lines, where two players alternately place black and white stones aiming to control the largest territory. The paradigm of Go artificial intelligence (AI) is revolutionized with the launch of DeepMind’s AlphaGo series (Silver et al., 2016; 2017; 2018; Schrittwieser et al., 2020), which leveraged the power of self-play to significantly elevate performance. In AlphaGo (Silver et al., 2016), the training regime can be split into three stages. In the first stage, supervised learning with expert data trains a fast policy network $p_{\pi}(a|s)$ for rollouts in MCTS and a precise policy network $p_{\sigma}(a|s)$. The second stage employs self-play combined with RL to get a refined policy network $p_{\rho}(a|s)$ based on $p_{\sigma}(a|s)$ and subsequently trains a value network $v_{\theta}(s)$. Specifically, $p_{\rho}(a|s)$ is refined by competing against a randomly selected historical version $p_{\rho^{-}}(a|s)$, similar to the MSS shown in Equ. (6), while $v_{\theta}(s)$ is trained using the game samples collected by self-play of $p_{\rho}(a|s)$. In the third stage, MCTS integrates the policy and value networks to select actions.
@@
Based on AlphaGo, AlphaGo Zero (Silver et al., 2017) does not require any expert data except game rules. It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities and the state value. Self-play is employed to generate data and refine $f_{\theta}(s)$ with the current best policy competing against itself, a process analogous to the MSS referenced in Equ. (5). A new policy to be incorporated into the policy pool must surpass a 55 percent win rate against its predecessor. AlphaZero (Silver et al., 2018) extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi. Concerning the self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network to generate samples by self-play without validation. Building upon AlphaZero, MuZero (Schrittwieser et al., 2020) takes learning from scratch to the next level, even operating without predefined game rules. The self-play process in MuZero operates similarly to that in AlphaZero. In practice, in addition to excelling in board games, MuZero also achieves state-of-the-art performance in Atari games.
4.1.2 Stratego
@@
4.2 Card Games
Unlike board games, card games are typically imperfect information games with a larger state space and greater complexity. Here, we introduce three representative card games: Texas Hold’em, DouDiZhu, and Mahjong, to illustrate how self-play is utilized in card games. Texas Hold’em is a repeated game rather than a stage game, so the focus is on the overall win-loss outcome rather than on individual rounds. Consequently, it is particularly well-suited for regret-minimization-based self-play methods, which aim to minimize overall regret rather than regret on a per-round basis. In contrast, DouDiZhu and Mahjong are typically regarded as stage games, so studies often employ self-play methods that do not rely on regret minimization.
4.2.1 Texas Hold’em
@@
Heads-up limit Texas Hold’em (HULHE), the simplest form with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus (Bowling et al., 2015). It utilizes fixed-point arithmetic with compression and regret-minimization-based self-play method CFR+ (Tammelin, 2014) to address the issues of storage and computation, respectively, resulting in superhuman performance. Deep CFR (Brown et al., 2019) combines regret-minimization-based self-play method CFR with deep neural networks. Furthermore, some studies do not adopt the regret-minimization-based series of self-play; instead, they use the traditional self-play method. NSFP (Heinrich & Silver, 2016) introduces a self-play method combined with end-to-end RL training, also achieving competitive performance in HULHE. Poker-CNN (Yakovenko et al., 2016) utilizes self-play to learn convolutional networks to solve a video version of HULHE.
After solving HULHE, research focus shifts to heads-up no-limit Texas Hold’em (HUNL) with significantly larger game states, approximately $10^{164}$ (Johanson, 2013). Thus, traversing the game tree as Cepheus does in HULHE is impossible. DeepStack (Moravčík et al., 2017) is a regret-minimization-based self-play method combined with continual re-solving. This method focuses on a subtree of limited depth and breadth and estimates the outcomes of the furthest reaches of this subtree. Libratus (Brown & Sandholm, 2018) develops a blueprint strategy, leveraging an enhanced version of regret-minimization-based self-play method MCCFR (Lanctot et al., 2009). Recursive Belief-based Learning (ReBeL) (Brown et al., 2020) introduces the public belief state (PBS), transforming the imperfect-information game into a perfect-information game with continuous state space. ReBeL utilizes the regret-minimization-based self-play method CFR-D, combined with RL and search, to train both the value and policy networks. Unlike the algorithms mentioned above, AlphaHoldem (Zhao et al., 2022a) does not use regret-minimization-based self-play. Instead, it proposes a $K$-best self-play method that always selects the top $K$ agents, based on their ELO ratings (Elo & Sloan, 1978), to generate samples.
Pluribus (Brown & Sandholm, 2019b), based on Libratus, addresses six-player no-limit Texas Hold’em. Similar to Libratus, Pluribus utilizes a regret-minimization-based self-play method MCCFR (Lanctot et al., 2009) to develop its blueprint strategy. It conducts a depth-limited search before executing actions. Different from Libratus, Pluribus maintains a streamlined policy pool of only four strategies, assuming that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently.
4.2.2 DouDiZhu

### v4

@@ 3.5.5. Estimation Variants of CFR
Although these CFR variants advance the field, they can’t directly solve large imperfect-information extensive-form games due to their reliance on tabular representations. The typical approach involves abstracting the original game, applying CFR to the abstracted game, and translating strategies back to the original. This abstraction is game-specific and relies heavily on domain knowledge. Additionally, smaller abstractions often yield suboptimal results. Given these challenges, Waugh et al. (2015) introduces Regression CFR (RCFR), which employs a shared regressor $\varphi(I,a)$ to estimate counterfactual regrets. Nevertheless, using regression trees as the regressor limits RCFR’s applicability to miniature games, and the necessity for manually crafted features remains a drawback. After advantage-based regret minimization (ARM) (Jin et al., 2018) merges CFR with deep RL in single-agent scenarios, a growing body of research has focused on applying CFR in conjunction with neural networks to multi-agent scenarios. Double Neural CFR (Li et al., 2020a) utilizes two neural networks: one for estimating counterfactual regrets and another for approximating the average strategy. In a similar vein, Deep CFR (Brown et al., 2019) leverages an advantage network $V(I,a|\theta_{p})$ to estimate counterfactual regrets with each player having a distinct hyperparameter $\theta_{p}$ and employs $\pi(I,a|\theta_{\pi})$ for strategy estimation after the training process of the advantage network. Since these two networks are trained in sequence rather than concurrently, the strategy for each intermediate iteration remains conditioned on the output of the advantage network: $h(k)=V(I,a|\theta_{p})$. Despite similarities, Deep CFR distinguishes itself from Double Neural CFR through its data collection and proven effectiveness in larger-scale poker games. Single Deep CFR (SD-CFR) (Steinberger, 2019) demonstrates training an average strategy network is unnecessary, with only an advantage network required. Building on the foundation of SD-CFR, DREAM (Steinberger et al., 2020) utilizes a learned baseline to maintain low variance in a model-free setting when only one action is sampled at each decision point. Moreover, advantage regret-matching actor-critic (ARMAC) (Gruslys et al., 2020) incorporates the thought of retrospective policy improvement.
@@
Table 1. Overview of representative self-play algorithms within our framework.
Algorithms MSS h(k) Categories $E$ Initialization of $\Pi$ Initialization of $\pi_{[k]}^{h}$
Vanilla Self-Play Equ. (5) $\emptyset$ Traditional Self-Play $E=1$ Lazy General
FP Equ. (6)
@@
DO NE-Based (Algo. 6) $\emptyset$ PSRO Series $E=1$ Lazy General
@@
FTW Equ. (10) $\emptyset$ Ongoing-Training-Based $E>1$ Immediate $\pi_{[k]}^{h}\leftarrow\pi_{[k]}(\cdot|h(k))$
NeuPL General $\sigma_{[k]}$
Simplex-NeuPL General $\sigma_{[k]}$
Vanilla CFR Equ. (5) $R^{k}(I,a)$ Regret-Minimization-Based $E=1$ Lazy $\pi_{[k]}^{h}\leftarrow\pi_{[k-1]}(\cdot|h(k-1))$
CFR+ $R^{+,k}(I,a)$
CFR-D $R^{k}(I_{*},a)$
@@ 3.6. Reassessment of the Framework
After introducing these four categories of self-play algorithms, we will further compare them in this section, explain the rationale behind categorizing the algorithms, and summarize the representative algorithms in Table 1. Moreover, we will illustrate the differences between our framework and other frameworks to demonstrate why our proposed framework is more general. Finally, we discuss how our framework can guide the development of future self-play algorithms by offering a unified perspective on their design and evaluation.
3.6.1. Categorization Rationale
@@
Traditional self-play algorithms and the PSRO series share many similarities. Initially, they require only one randomly initiated policy, and the policy population expands as training progresses. Therefore, in our framework, we use lazy initialization to initialize the policy population and set $E=1$ for these two categories. The interaction matrix is typically a lower triangular matrix (Corollary 3.1 and Corollary 3.2). The primary difference between the PSRO series and traditional self-play algorithms is that the PSRO series employs more complex MSSs to handle tasks with intricate game-theoretical requirements. For example, $\alpha$-PSRO (Muller et al., 2020) specifically utilizes an $\alpha$-rank-based MSS to tackle multi-player general-sum games. In other tasks, traditional self-play algorithms are more commonly used to reduce the computational cost associated with complex MSSs in the PSRO series.
@@
Unlike the two previously mentioned categories, the ongoing-training-based series adopts a different paradigm. Instead of gradually expanding the policy population and relying on newer policies to be stronger, this approach strengthens all policies simultaneously at each epoch. This method alleviates issues such as early truncation and repeated skill relearning that occur in the above two categories. To integrate this category into our framework, immediate initialization is used for the policy population, and $\pi_{[k]}(\cdot|h(k))$ is utilized to initialize $\pi_{[k]}^{h}$ to ensure that policy updates are self-referential. Also, the interaction matrix is generally not a lower triangular matrix (Corollary 3.3).
Lastly, the regret-minimization-based series follows an online learning paradigm that focuses on optimizing overall performance across time rather than on individual episodes. Representative methods such as CFR and its variants are specifically designed for extensive-form imperfect information games, making them particularly effective in such settings. The core mechanism of regret-minimization-based training is to iteratively update the regrets associated with different strategies. However, tracking and accumulating regret for every information set incurs substantial memory overhead, especially in large game trees. Furthermore, available theoretical guarantees of convergence to a Nash equilibrium have so far been established only for two-player zero-sum games; achieving comparable assurances in multiplayer or general-sum settings remains an open research problem. Our framework uses $h(k)$ to store this information. Since the policies are determined by $h(k)$, only the most recent policy is relevant. Therefore, the interaction matrix is a unit lower shift matrix (Corollary 3.4). We also do not need to actually initialize the whole policy population and only need to use $\pi_{[k-1]}(\cdot|h(k-1))$ to initialize $\pi_{[k]}^{h}$ in the training process.
3.6.2. Comparing with Existing Frameworks
@@
Our framework is built upon PSRO (Lanctot et al., 2017) and NeuPL (Liu et al., 2022c). Here, we outline the differences between our framework and these existing frameworks. The primary distinction between our framework and PSRO is the use of an interaction matrix $\Sigma:=\{\sigma_{[k]}\}_{k=1}^{K}\in\mathbb{R}^{K\times C_{1}}$ to represent the opponent sampling strategy, allowing for the integration of more complex competitive dynamics. Moreover, in our framework, $\sigma_{[k]}$ denotes the opponent sampling strategy, which specifies how to sample opponents’ policies against policy $k$ rather than being the meta-strategy of policy $k$. Additionally, our framework incorporates a policy condition function $h(k)$, making it more general than NeuPL, where policies are conditioned on $\sigma_{[k]}$. This enhancement gives our framework greater expressiveness. Furthermore, we describe how to compute the oracle (Line 5 in Alg. 1) in three different ways (Alg. 2, Alg. 3 and Alg. 4) to provide a clearer understanding. Also, to the best of our knowledge, our framework is the first self-play framework to integrate the regret-minimization-based series, which is a significant self-play paradigm.
3.6.3. Implications for Future Self-Play Algorithm Design
Beyond merely organizing existing algorithms, our framework provides insights into the design of future self-play methods. First, the four algorithmic categories we identify are not isolated; mechanisms developed in one family can be transplanted into another. Second, although most previous work has concentrated on the ORACLE and MSS, our unified framework uncovers additional, under-explored dimensions that warrant further investigation. The remainder of this section first illustrates how techniques can migrate across categories and then presents two specific research directions that arise when the full framework design space is considered.
Cross-category inspiration. Recent advances in self-play have explicitly blended techniques across different algorithmic categories, yielding hybrid methods that transcend traditional boundaries. RM-BR (Johanson et al., 2012a) pairs RM for one player with a BR oracle for the other, leveraging RM’s equilibrium guarantees and exploitative strength. Anytime PSRO (McAleer et al., 2022b) proposes RM-BR DO by incorporating a regret-minimization component into DO to ensure that exploitability monotonically decreases as new strategies are added. ODO (Dinh et al., 2022) fuses another online no-regret learning method multiplicative weights update (MWU) (Freund and Schapire, 1999) with DO. Our framework naturally subsumes ODO. Policies are still added iteratively, and the auxiliary information $h(k)$ records the average regret-based losses. MWU acts as the ORACLE. Crucially, unlike the original DO, this MWU oracle need not compute a full Nash equilibrium, reducing computational cost; yet ODO still provably converges to a Nash equilibrium in two-player zero-sum normal-form games while preserving rationality, allowing agents to exploit sub-optimal opponents. Taken together, these studies show that innovation thrives when regret-minimization-based series intersect with PSRO series, and they suggest that integrating other categories could be just as fruitful.
Prioritised re-training when $E\gg 1$. When each policy is reused across many encounters ($E\gg 1$), repeatedly updating all members becomes computationally expensive. A natural open question is: which policies are worth further improvement? For instance, we can formalize this as a multi-armed bandit: give every policy a score that estimates how much new information an additional update would add (e.g., its expected information gain), and let that score decide where to spend oracle calls. This bandit scheduling automatically balances exploitation (further refining the strongest policies) with exploration (re-examining under-trained or atypical policies that might expose blind spots).
Auxiliary information $h(k)$ and knowledge transfer. The history kernel $h(k)$ need not serve merely as a passive log of past interactions; instead, it can be designed to encode rich auxiliary domain knowledge, such as exploitable behavioral patterns, partial payoff structures, or latent embeddings of opponents. By treating $h(k)$ as a learnable memory object that can be inherited or shared across subsets of agents, we enable a form of continual-learning self-play. In this setting, knowledge acquired by one agent can propagate throughout the population organically, without requiring intervention from an external oracle. This facilitates more efficient collective learning and strategic evolution within multi-agent systems.
4. Empirical Analysis
@@ 4.1. Board Games
Board games, the majority of which are perfect information games, were previously revolutionized by the introduction of two essential techniques: position evaluation and MCTS (Coulom, 2006; Kocsis and Szepesvári, 2006). These methodologies, with minor modifications, demonstrated superhuman effectiveness in solving board games such as chess (Campbell et al., 2002), checkers (Schaeffer et al., 1992), othello (Buro, 1999), backgammon (Tesauro and Galperin, 1996), and Scrabble (Sheppard, 2002). In contrast, the application of these techniques to the game of Go with an estimated $2.1\times 10^{170}$ legal board configurations, only enabled performance at the amateur level (Bouzy and Helmstetter, 2004; Coulom, 2007; Baudiš and Gailly, 2011; Enzenberger et al., 2010; Gelly and Silver, 2007). In light of this, our discussion will specifically focus on the game of Go to illustrate the application of vanilla self-play. In addition to Go, we will explore Stratego, a board game with imperfect information, in contrast to most board games that only involve perfect information.
4.1.1. Go
@@
Go is an ancient board game, which is played on a grid of 19x19 lines, where two players alternately place black and white stones aiming to control the largest territory. The paradigm of Go artificial intelligence (AI) is revolutionized with the launch of DeepMind’s AlphaGo series (Silver et al., 2016, 2017, 2018; Schrittwieser et al., 2020), which leveraged the power of vanilla self-play to significantly elevate performance. In AlphaGo (Silver et al., 2016), the training regime can be split into three stages. In the first stage, supervised learning with expert data trains a fast policy network $p_{\pi}(a|s)$ for rollouts in MCTS and a precise policy network $p_{\sigma}(a|s)$. The second stage employs vanilla self-play combined with RL to get a refined policy network $p_{\rho}(a|s)$ based on $p_{\sigma}(a|s)$ and subsequently trains a value network $v_{\theta}(s)$. Specifically, $p_{\rho}(a|s)$ is refined by competing against a randomly selected historical version $p_{\rho^{-}}(a|s)$, similar to the MSS shown in Equ. (6), while $v_{\theta}(s)$ is trained using the game samples collected by vanilla self-play of $p_{\rho}(a|s)$. In the third stage, MCTS integrates the policy and value networks to select actions.
@@
Based on AlphaGo, AlphaGo Zero (Silver et al., 2017) does not require any expert data except game rules. It utilizes only one network $f_{\theta}(s)$ to concurrently predict the action probabilities $\text{Pr}(a|s)$ and the state value $v$. Each move is generated by MCTS with guidance from $f_{\theta}(s)$ to aid MCTS expansion, unlike the rollouts used in AlphaGo. Vanilla self-play is employed to generate data and refine $f_{\theta}(s)$ with the current best policy competing against itself, a process analogous to the MSS referenced in Equ. (5). A new policy to be incorporated into the policy pool must surpass a 55 percent win rate against its predecessor, aligning with the stipulation set in Algo. 2 at Line 1. AlphaZero (Silver et al., 2018) extends AlphaGo Zero to include games beyond Go, such as Chess and Shogi. Notably, a draw is introduced as an additional expected outcome, and data augmentation is omitted due to the asymmetry of Chess and Shogi. Concerning the vanilla self-play procedure, the only difference between AlphaZero and AlphaGo Zero is that AlphaZero utilizes the newly updated network without the validation process present in AlphaGo Zero. Building upon AlphaZero, MuZero (Schrittwieser et al., 2020) takes learning from scratch to the next level, even operating without predefined game rules. MuZero incorporates ideas from model-based RL to model the dynamics of games. In addition to the prediction network $f$ (similar to the networks in AlphaGo Zero and AlphaZero), MuZero introduces a dynamics network $g$ to model the MDP and a representation network $h$ to map observations to hidden states. These three networks are trained jointly. MuZero employs MCTS guided by the three networks above to make decisions. The vanilla self-play in MuZero operates similarly to that in AlphaZero. In practice, in addition to excelling in board games like Go, MuZero also achieves state-of-the-art performance in Atari games.
4.1.2. Stratego
@@
4.2. Card Games
Unlike board games, card games are typically imperfect information games with a larger state space and greater complexity. Some regret-minimization-based self-play methods, particularly CFR and its variants, are well-suited for handling imperfect information and have achieved strong results in card games. However, recent studies have shown that other self-play methods without regret minimization can also perform well in such settings. We introduce three representative card games: Texas Hold’em, DouDiZhu, and Mahjong, to illustrate how self-play is utilized in card games.
4.2.1. Texas Hold’em
@@
Heads-up limit Texas Hold’em (HULHE), the simplest form with approximately $3.16\times 10^{17}$ game states, was not solved until the introduction of Cepheus (Bowling et al., 2015). It utilizes fixed-point arithmetic with compression and regret-minimization-based self-play method CFR+ (Tammelin, 2014) to address the issues of storage and computation, respectively, resulting in superhuman performance. Deep CFR (Brown et al., 2019) combines regret-minimization-based self-play method CFR with deep neural networks. Furthermore, some studies do not adopt the regret-minimization-based series of self-play; instead, they use the traditional self-play method. NSFP (Heinrich and Silver, 2016) introduces a self-play method combined with end-to-end RL training, also achieving competitive performance in HULHE. Poker-CNN (Yakovenko et al., 2016) utilizes self-play to learn convolutional networks to solve a video version of HULHE.
After solving HULHE, research focus shifts to heads-up no-limit Texas Hold’em (HUNL) with significantly larger game states, approximately $10^{164}$ (Johanson, 2013). Thus, traversing the game tree as Cepheus does in HULHE is impossible. DeepStack (Moravčík et al., 2017) is a regret-minimization-based self-play method combined with continual re-solving. This method focuses on a subtree of limited depth and breadth and estimates the outcomes of the furthest reaches of this subtree. Libratus (Brown and Sandholm, 2018) develops a blueprint strategy, leveraging an enhanced version of regret-minimization-based self-play method MCCFR (Lanctot et al., 2009). Recursive Belief-based Learning (ReBeL) (Brown et al., 2020) introduces the public belief state (PBS), transforming the imperfect-information game into a perfect-information game with continuous state space. ReBeL utilizes the regret-minimization-based self-play method CFR-D, combined with RL and search, to train both the value and policy networks. Unlike the algorithms mentioned above, AlphaHoldem (Zhao et al., 2022a) does not use regret-minimization-based self-play. Instead, it proposes $K$-best self-play that selects the top $K$ policies, based on their ELO ratings (Elo and Sloan, 1978), to generate samples.
Pluribus (Brown and Sandholm, 2019b), based on Libratus, addresses six-player no-limit Texas Hold’em. Similar to Libratus, Pluribus utilizes a regret-minimization-based self-play method MCCFR (Lanctot et al., 2009) to develop its blueprint strategy. It conducts a depth-limited search before executing actions. Different from Libratus, Pluribus maintains a streamlined policy pool of only four strategies, assuming that its opponents might adjust their strategies during gameplay among these four strategies, which allows Pluribus to manage complexity more efficiently.
4.2.2. DouDiZhu